HIGH RESOLUTION LASER SPECTROSCOPY OF RHODIUM MONOPHOSPHIDE

\underline{A. G. ADAM}, L. E. DOWNIE, and M. E. SLANEY, Chemistry Department, and Centre for Lasers, and Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, NB, E3B 6E2; and W. J. BALFOUR and R. LI, Chemistry Department, University of Victoria, Victoria, BC, V8W 3P6.

High resolution laser spectra of RhP have been acquired in the visible region of the spectrum. The molecules were produced by laser ablation of a rhodium target rod followed by reaction with 2% PH_{3} seeded in He in a pulsed supersonic jet. At least three electronic transitions with extensive upper state vibrational progressions have been observed by laser-induced fluorescence between 420 nm and 510 nm. All the bands are degraded to longer wavelengths and have an R branch which rapidly forms a head and a more open P branch. First lines in the bands were seen allowing line assignments which are consistent with ^{1}Σ - ^{1}Σ transitions. Dispersed fluorescence from the bands give a ground state vibrational interval of approximately 507 cm^{−1} plus also indicate the presence of low-lying electronic states at 2454 cm^{−1} and 4691 cm^{−1}. Molecular constants for the excited and ground states will be presented.